# Modeling, Inference, and Prediction in Mobility-Based Compartmental Models for Epidemiology

### v1

@@ 2.4 Mean mobility
Mobility plays a crucial role in the mobility-based compartmental models. We examine the mean mobility of three populations. We focus on the mobility-based SIR model, which is defined by the following equations
$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},$ (17a)
$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},-\gamma I(m,t)$ (17b)
$\frac{\partial}{\partial t}R(m,t)$ $=\gamma I(m,t).$ (17c)
We define $Q_{k}^{\varphi}$ as the $k$th moment with respect to a nonnegative function $\varphi(m,t)$, which is
@@
Let ${Q_{k}^{S}}(t)$ be the $k$th moment of $S(m,t)$ which is the solution of the mobility-based SIR model (17). If $f$ is not a Dirac delta distribution, then for any $k\geq 0$, the ratio ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases with respect to $t$.
@@
We multiply $m^{k}$ to (17a) and integrate $m$ over $[0,1]$. We obtain
@@
By using (19), we obtain the time derivative of ${Q_{k+1}^{S}}/{Q_{k}^{S}}$
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{k+1}^{S}}{Q_{k}^{S}}\right)=\frac{1}{Q_{k}^{S}Q_{k}^{S}}\left(Q_{k}^{S}\dot{Q}_{k+1}^{S}-\dot{Q}_{k}^{S}Q_{k+1}^{S}\right)=-\frac{\beta Q_{1}^{I}}{Q_{k}^{S}Q_{k}^{S}}\left(Q_{k}^{S}Q_{k+2}^{S}-Q_{k+1}^{S}Q_{k+1}^{S}\right).$ (20)
From Hölder’s inequality, we know that
$Q_{k}^{S}Q_{k+2}^{S}-Q_{k+1}^{S}Q_{k+1}^{S}>0,$ (21)
which implies that the time derivative in (20) is negative, so that ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases. ∎
For $k=0$, the ratio in Theorem 1 simplifies to $Q_{1}^{S}(t)/Q_{0}^{S}(t)$, representing the mean mobility of the susceptible population. Thus, in the mobility-based SIR model (17), the mean mobility of the susceptible population decreases over time $t$. However, this monotonic property does not necessarily apply to the infected and recovered populations, as illustrated in Figure 4. In addition, if $f$ is a Dirac delta distribution, then the mean mobility remains constant and is independent of time $t$.
@@
[…]
@@ 2.5 Final pandemic size
The final pandemic size refers to the total number of individuals infected throughout an epidemic and is a crucial quantity to assess disease transmission dynamics. For the classical SIR model with a basic reproduction number $\mathcal{R}_{0}$, it is well known that the final pandemic size $R_{\infty}$ satisfies the equation
@@
if the initial infected population $I\approx 0$. In the real world, the population exhibits heterogeneous mobility, with high-mobility individuals likely being infected first during the early pandemic stages. Once recovered, they provide some protection to low-mobility individuals. In practice, $\mathcal{R}_{0}$ is typically estimated from the initial exponential growth rate, which is biased toward the infection rate of high-mobility individuals, thereby resulting in an overestimation of the final pandemic size. In this section, we derive the final pandemic size of the mobility-based SIR model (17) with the proportional initial condition (6). We rigorously demonstrate why the classic SIR model tends to overestimate the final size of pandemics, assuming the same $\mathcal{R}_{0}$ as in the mobility-based SIR model.
@@
Let $R_{\infty}$ be the final pandemic size of the mobility-based SIR model (17) with the proportial initial condition (6). Assume $I=0\approx 0$, then $R_{\infty}$ satisfies
$R_{\infty}+\langle f\exp\left(-\alpha m\beta/\gamma\right)\rangle$ $=1,$ (23)
where $\alpha$ is the solution of $\alpha+\langle mf\exp\left(-\alpha m\beta/\gamma\right)\rangle=\langle mf\rangle$ on the interval $[0,1]$.
@@
By solving (2) for $S(m,t)$, we obtain that
@@
We use the initial condition (6) and assume that $S(m,0)\approx f(m)$, then $\alpha=\langle mR\rangle(\infty)$ satisfies
$\alpha+\langle mf\exp\left(-\alpha m\beta/\gamma\right)\rangle=\langle mf\rangle.$ (27)
Integrating (24) and taking $t$ to infinite, we obtain that $R_{\infty}=\langle R\rangle(\infty)$ satisfies (23). ∎
Theorem 2 offers a method to compute the final pandemic size without the need to simulate the model (17) over time. The process involves first determining $\alpha$ by finding the root of (27) within the interval $[0,1]$. Once $\alpha$ is obtained, we then solve (23) to find $R_{\infty}$. Additionally, once the parameters $\beta$ and $\gamma$ are fixed, the mobility distribution $f(m)$ solely determines both the final pandemic size $R_{\infty}$ and the basic reproduction number $\mathcal{R}_{0}$ as given in (14). We will demonstrate in Theorem 3 that among all distributions that yield the same basic reproduction number, the Dirac delta distribution $f_{\delta}$, which corresponds to a homogeneous mobility of the population, maximizes the final pandemic size.
@@
Let $R_{\infty}$ and $R^{\delta}_{\infty}$ be the final pandemic sizes associated with a distribution $f$ and a Dirac delta distribution $f_{\delta}$, respectively. Suppose that $f$ and $f_{\delta}$ yield the same basic reproduction number $\mathcal{R}_{0}$ (14), it holds that $R_{\infty}\leq R^{\delta}_{\infty}$ with the equality occurring only if $f=f_{\delta}$.
Theorem 2 provides an approach to compute the final pandemic size. We introduce another perspective to represent the final pandemic size to prove Theorem 3.
@@
From (17), we have
$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}t}=-\beta Q_{1}^{S}Q_{1}^{I},\quad\frac{\mathrm{d}Q_{0}^{R}}{\mathrm{d}t}=\gamma Q_{0}^{I},\quad t\in[0,\infty).$ (28)
Since $Q_{0}^{R}(t)$ monotonely increases with respect to $t$, so we have a one-to-one correspondence between $Q_{0}^{R}(t)$ and $t$. Therefore, we can view all functions of $t$ (including $Q_{0}^{S}$) as a function of $Q_{0}^{R}$. Equation (28) becomes
$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}=-\frac{\beta}{\gamma}\widetilde{H}(Q_{0}^{R})Q_{0}^{S},\quad Q_{0}^{R}\in[Q_{0}^{R}(0),Q_{0}^{R}(\infty)),$ (29)
where
$\widetilde{H}(Q_{0}^{R})=H(t)=\frac{Q_{1}^{S}(t)Q_{1}^{I}(t)}{Q_{0}^{S}(t)Q_{0}^{I}(t)}.$ (30)
Recall that $Q_{0}^{R}(0)=0$. We integrate (29) over the interval $[0,Q_{0}^{R}(\infty)]$ and obtain
$Q_{0}^{S}(\infty)=\exp\left(-\frac{\beta}{\gamma}\int_{0}^{Q_{0}^{R}(\infty)}\widetilde{H}(\alpha)\leavevmode\nobreak\ \mathrm{d}\alpha\right).$ (31)
Using $\lim_{t\to\infty}Q_{0}^{I}(t)=0$ and change of variable, we obtain
$Q_{0}^{R}(\infty)+\exp\left(-\frac{\beta}{\gamma}\int_{0}^{\infty}H(t)\leavevmode\nobreak\ \mathrm{d}Q_{0}^{R}(t)\right)=1.$ (32)
As $R_{\infty}^{\delta}$ satisfies
$R^{\delta}_{\infty}+\exp\left(-\mathcal{R}_{0}R^{\delta}_{\infty}\right)=1,$ (33)
it is sufficient to prove that
$H(t)\leq\frac{\gamma}{\beta}\mathcal{R}_{0}=\langle m^{2}f\rangle,\quad\text{for all\leavevmode\nobreak\ }t>0,$ (34)
where the equal sign holds only if $f=f_{\delta}$. It is not difficult to verify that the equality holds when $f=f_{\delta}$. We will assume that $f\neq f_{\delta}$ and prove that $H(t)<\langle m^{2}f\rangle$ in the following proof.
From Theorem 1 for $k=1$ and Lemma 4, we have
$\frac{Q_{1}^{S}(t)}{Q_{0}^{S}(t)}<\langle mf\rangle\quad\text{and}\quad\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}\leq\frac{\langle m^{2}f\rangle}{\langle mf\rangle},\quad\text{for all\leavevmode\nobreak\ }t>0.$ (35)
Combining it with (30), we complete the proof. ∎
Lemma 4.
Let $I(m,t)$ be the solution of the mobility-based SIR model (17) with the proportional initial condition (6). For all $t>0$, it holds that
with equality occurring only if $f$ is a Dirac delta distribution.
Proof.
If $f$ is a Dirac delta distribution $f(m)=\delta(m-m_{0})$, then both sides of (36) equal $m_{0}$. We assume that $f$ is not a Dirac delta distribution in the following proof.
@@
We denote the instants that $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}\right)$ changes its sign as $t_{k}$ and partition the time interval $(0,\infty)$ into open intervals $\Omega_{k}$ with $t_{k}$ as indicated in Figure 5.
@@
which implies that $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}}/{Q_{0}^{I}}\right)$ is nonnegative on the interval $\Omega_{1}$. Due to the smoothness, the signs of (37) interlace on $\Omega_{1}$, $\Omega_{2}$, $\Omega_{3}$, $\cdots$.
If $t\in\cup_{k=1}^{\infty}\overline{\Omega_{2k-1}}$, then we know (37) is nonnegative, which implies that
@@
where the second inequality uses Theorem 1 for the case $k=1$. If $t\in\cup_{k=1}^{\infty}\Omega_{2k}$ (supposing $t\in\Omega_{2k}$), then ${Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}$ decreases with respect to $t$ on the interval $\Omega_{2k}$. Due to continuity, we have
@@
where the second inequality holds from (39).
@@ 3 Learning the mobility distribution from data
Mobility holds critical significance in epidemic models as it directly influences the spread and dynamics of infectious diseases within populations. Understanding how people move between locations allows for more accurate predictions of disease transmission and enhances the precision of epidemic models to reflect real-world scenarios. However, the mobility of a population involves complex factors, including travel patterns, migration, and commuting, which are often impossible to measure directly in real-world situations. Realistically, the available data in epidemics is the number of infected individuals, which translates to the ratio of infected agents at time $t$:
@@
in the mobility-based SIRS model (4). In this section, we infer the mobility distribution $f(m)\in\mathcal{L}^{1}(0,1)$ from the time series of $Q_{0}^{I}(t)$, offering both a theoretical guarantee of unique inference and a machine-learning-based numerical approach.
3.1 Theoretical guarantees

### v2

@@ 2.4 Mean mobility
Mobility plays a crucial role in the mobility-based compartmental models. We focus on the mobility-based SIR model, which is defined by the following equations,
$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}m\bar{m}S(m,t)I(\bar{m},t)\leavevmode\nobreak\ \mathrm{d}\bar{m},$ (16a)
$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}m\bar{m}S(m,t)I(\bar{m},t)\leavevmode\nobreak\ \mathrm{d}\bar{m}-\gamma I(m,t),$ (16b)
$\frac{\partial}{\partial t}R(m,t)$ $=\gamma I(m,t),$ (16c)
and examine the mean mobility of the susceptible compartment. Let $Q_{k}^{\varphi}$ be the $k$th moment with respect to a nonnegative function $\varphi(m,t)$, which means
@@
Under this definition, the ratio of the susceptible population is $Q_{0}^{S}(t)$, and the mean mobility of the susceptible compartment is $Q_{1}^{S}(t)/Q_{0}^{S}(t)$. We show the monotone property associated with the susceptible population in Theorem 2.
@@
Let ${Q_{k}^{S}}(t)$ be the $k$th moment of $S(m,t)$ in the mobility-based SIR model (16). If $f$ is not a Dirac delta distribution, then for all $k\geq 0$, the ratio ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases with respect to time $t$.
@@
By multiplying $m^{k}$ to (16a) and integrating $m$ over $[0,1]$, we obtain that
@@
Using (18), we obtain the time derivative of ${Q_{k+1}^{S}}/{Q_{k}^{S}}$
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{Q_{k+1}^{S}}{Q_{k}^{S}}\right)=\frac{1}{(Q_{k}^{S})^{2}}\left(Q_{k}^{S}\dot{Q}_{k+1}^{S}-\dot{Q}_{k}^{S}Q_{k+1}^{S}\right)=-\frac{\beta Q_{1}^{I}}{(Q_{k}^{S})^{2}}\left[Q_{k}^{S}Q_{k+2}^{S}-(Q_{k+1}^{S})^{2}\right].$ (19)
From Hölder’s inequality, we have that
$Q_{k}^{S}Q_{k+2}^{S}-(Q_{k+1}^{S})^{2}>0,$ (20)
which implies that the time derivative in (19) is negative, so that ${Q_{k+1}^{S}}(t)/{Q_{k}^{S}}(t)$ strictly decreases over time $t$. ∎
For $k=0$, the ratio in Theorem 2 simplifies to $Q_{1}^{S}(t)/Q_{0}^{S}(t)$, representing the mean mobility of the susceptible population. Thus, in the mobility-based SIR model (16), the mean mobility of the susceptible population decreases over time $t$. However, this monotonic property does not necessarily apply to the infected and recovered populations, as illustrated in Figure 4. In addition, if $f$ is a Dirac delta distribution, then the entire population has a constant mobility, causing the mean mobility to remain unchanged over time.
@@
Figure 4: Time evolution of the mean mobility of three compartments in the mobility-based SIR model (16) with parameters $\beta=1.5$, $\gamma=0.13$, $I_{0}=0.0001$, and $f=f_{1}$ in (7). The mean mobility of the susceptible population decreases strictly over time, while the mean mobility of the infected and the recovered populations increases first and then decreases as time progresses.
@@ 2.5 Final pandemic size
The final pandemic size, defined as the total number of individuals infected over the course of an epidemic, is a critical quantity for assessing disease transmissions. For the classical SIR model, it is well known that the final pandemic size $R_{\infty}$ satisfies
@@
where $\mathcal{R}_{0}$ is the basic reproduction number. Equation (21) often overestimates the final pandemic size. For instance, when $\mathcal{R}_{0}=2$, Equation (21) suggests that $R_{\infty}$ is around $0.8$, implying that nearly $80\%$ of the population will ultimately be infected, which is a scenario rarely observed in practice.
In reality, populations have heterogeneous mobility. High-mobility individuals are more likely to be infected early in a pandemic, and once they recover, they provide indirect protection to low-mobility individuals. Typically, $\mathcal{R}_{0}$ is estimated from the initial exponential growth rate, which is skewed by the higher infection rates among high-mobility individuals. This bias leads to an overestimation of the final pandemic size. Take the recent COVID-19 pandemic as an example. Studies estimate the basic reproduction number $\mathcal{R}_{0}$ for COVID-19 to be between 2 and 5 [16, 29, 6] in early 2020. However, later variants, such as Delta and Omicron, prove significantly more contagious. Studies suggest that the $\mathcal{R}_{0}$ of the Omicron variant is as high as 10 [27, 19, 2], which implies, according to the final pandemic size estimation in (21), that nearly the entire population would eventually become infected. However, empirical data indicate otherwise. As of December 2021, after two years of multiple infection waves caused by various SARS-CoV-2 variants, the seroprevalence in the United States is only 33.5%. Moreover, the largest infection wave, caused by the Omicron variant from December 2021 to February 2022, infects less than one-quarter of the total population [4]. These numbers are far below the final pandemic size suggested by the classical SIR model.
In this section, we derive the final pandemic size of the mobility-based SIR model (16) with the proportional initial condition (6). We provide a quantitative explanation of why the classical SIR model tends to overestimate the final pandemic size in comparison to our mobility-based SIR model (16).
@@
Let $R_{\infty}$ be the final pandemic size of the mobility-based SIR model (16) with the proportial initial condition (6). Then the final pandemic size $R_{\infty}$ satisfies
$R_{\infty}+\langle f\exp\left(-\frac{\alpha m\beta}{\gamma}\right)\rangle$ $=1,$ (22)
where $\alpha$ is the root on the interval $[0,1]$ of the equation
$\alpha+\langle mf\exp\left(-\frac{\alpha m\beta}{\gamma}\right)\rangle=\langle mf\rangle.$ (23)
@@
We solve (16) for $S(m,t)$ and obtain that
@@
Using the initial condition (6) and $S(m,0)\approx f(m)$, we find that $\alpha=\langle mR\rangle(\infty)$ satisfies
$\alpha+\langle mf\exp\left(-\frac{\alpha m\beta}{\gamma}\right)\rangle=\langle mf\rangle.$ (27)
We integrate (24) over $m$ and take $t$ to infinite, then we obtain that $R_{\infty}=\langle R\rangle(\infty)$ satisfies (22). ∎
Theorem 3 offers a method to compute the final pandemic size of the model (16) efficiently. We first determine $\alpha$ by finding the root of (27) within the interval $[0,1]$ and then use the value of $\alpha$ to compute $R_{\infty}$ using (22). Once the parameters $\beta$ and $\gamma$ are fixed, the mobility distribution $f(m)$ solely determines both the final pandemic size $R_{\infty}$ and the basic reproduction number $\mathcal{R}_{0}$ as given in (12). In Theorem 4, we prove that among all distributions that yield the same basic reproduction number $\mathcal{R}_{0}$, the Dirac delta distribution $f_{\delta}$, which corresponds to a population with the same mobility, maximizes the final pandemic size.
@@
Let $R_{\infty}$ and $R^{\delta}_{\infty}$ be the final pandemic sizes associated with the mobility distribution $f$ and the Dirac delta mobility distribution $f_{\delta}$, respectively. Suppose that $f$ and $f_{\delta}$ yield the same basic reproduction number $\mathcal{R}_{0}$ (12), then $R_{\infty}\leq R^{\delta}_{\infty}$ and the equality occurs only if $f=f_{\delta}$.
Theorem 4 elucidates why the classical SIR model tends to overestimate the final pandemic size compared to our mobility-based model. The classical SIR model assumes a homogeneous population, which corresponds to a Dirac delta mobility distribution $f_{\delta}$ in the framework of our model (16). Its final pandemic size $R_{\infty}^{\delta}$ is greater than that of any other mobility distribution, given that all distributions share the same basic reproduction number $\mathcal{R}_{0}$.
We prepare the proof of Theorem 4 with the following lemma.
Lemma 5.
Let $I(m,t)$ be the mobility distribution of the infected population in the mobility-based SIR model (16) with the proportional initial condition (6). Then for all $t\geq 0$, it holds that
and the equality occurs only if $f$ is a Dirac delta distribution.
@@
If $f$ is a Dirac delta distribution $f(m)=\delta(m-m_{0})$, then both sides of (28) equal $m_{0}$. We assume that $f$ is not a Dirac delta distribution in the rest of the proof.
@@
We denote the time that $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}\right)$ changes its sign as $t_{k}$ and partition the time interval $(0,\infty)$ into open intervals $\Omega_{k}$ as shown in Figure 5.
@@
so $\frac{\mathrm{d}}{\mathrm{d}t}\left({Q_{1}^{I}}/{Q_{0}^{I}}\right)$ is positive on the interval $\Omega_{1}$. Due to the smoothness, the signs of (29) interlace on $\Omega_{1}$, $\Omega_{2}$, $\Omega_{3}$, $\ldots$.
If $t\in\cup_{k=1}^{\infty}{\Omega}_{2k-1}$ or $t\in\{t_{1},t_{2},\ldots\}$, then we know (29) is nonnegative, which implies that
@@
where the second inequality uses Theorem 2 for the case $k=1$. If $t\in\cup_{k=1}^{\infty}\Omega_{2k}$ (supposing $t\in\Omega_{2k}$), then ${Q_{1}^{I}(t)}/{Q_{0}^{I}(t)}$ decreases strictly with respect to $t$ on the interval $\Omega_{2k}$. Due to continuity, we have
@@
where the second inequality holds because of (31). Equations (31) and (32) complete the proof. ∎
Now, we turn to the proof of Theorem 4.
Proof.
From (16), we have
$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}t}=-\beta Q_{1}^{S}Q_{1}^{I},\quad\frac{\mathrm{d}Q_{0}^{R}}{\mathrm{d}t}=\gamma Q_{0}^{I},\quad\text{for all\leavevmode\nobreak\ }t\geq 0.$ (33)
Since $Q_{0}^{R}(t)$ increases strictly with respect to $t$, we have a one-to-one correspondence between $Q_{0}^{R}$ and $t$. Using the chain rule, we have
$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}\frac{\mathrm{d}Q_{0}^{R}}{\mathrm{d}t}=-\beta Q_{1}^{S}Q_{1}^{I},$ (34)
yielding
$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}=-\frac{\beta}{\gamma}\frac{Q_{1}^{S}Q_{1}^{I}}{Q_{0}^{I}}.$ (35)
Due to the one-to-one correspondence between $t$ and $Q_{0}^{R}$, we can view all functions of $t$ (including $Q_{0}^{S}$) as a function of $Q_{0}^{R}$ and rewrite (35) as
$\frac{\mathrm{d}Q_{0}^{S}}{\mathrm{d}Q_{0}^{R}}=-\frac{\beta}{\gamma}\widetilde{H}(Q_{0}^{R})Q_{0}^{S},$ (36)
where $\widetilde{H}$ is a function of $Q_{0}^{R}$ and defined by
$\widetilde{H}(Q_{0}^{R}(t))=\frac{Q_{1}^{S}(t)Q_{1}^{I}(t)}{Q_{0}^{S}(t)Q_{0}^{I}(t)}.$ (37)
Recall that $Q_{0}^{R}(0)=0$ and $Q_{0}^{R}(\infty)=R_{\infty}$. We integrate (36) over the interval $[0,R_{\infty}]$ and use $Q_{0}^{S}(\infty)=1-R_{\infty}$, obtaining
$R_{\infty}+\exp\left(-\frac{\beta}{\gamma}\int_{0}^{R_{\infty}}\widetilde{H}(\alpha)\leavevmode\nobreak\ \mathrm{d}\alpha\right)=1.$ (38)
Recall that $R_{\infty}^{\delta}$ satisfies (21). We rewrite (21) as
$R^{\delta}_{\infty}+\exp\left(-\int_{0}^{R_{\infty}^{\delta}}\mathcal{R}_{0}\leavevmode\nobreak\ \mathrm{d}\alpha\right)=1.$ (39)
From Theorem 2 for $k=1$ and Lemma 5, we have
$\frac{Q_{1}^{S}(t)}{Q_{0}^{S}(t)}<\langle mf\rangle\quad\text{and}\quad\frac{Q_{1}^{I}(t)}{Q_{0}^{I}(t)}<\frac{\langle m^{2}f\rangle}{\langle mf\rangle},\quad\text{for all\leavevmode\nobreak\ }t>0.$ (40)
Using (12), (37), and (40), we obtain that
$\frac{\beta}{\gamma}\widetilde{H}(\alpha)\leq\mathcal{R}_{0},\quad\text{for all\leavevmode\nobreak\ }\alpha\in[0,R_{\infty}),$ (41)
and the equality holds only if $f$ is a Dirac delta distribution (i.e., $f=f_{\delta}$).
From (38),(39), and (41), we obtain that $R_{\infty}\leq R_{\infty}^{\delta}$ and the equality holds only if $f=f_{\delta}$. ∎
@@ 3 Learning the mobility distribution from data
The mobility distribution plays a significant role in the mobility-based compartmental models. A realistic mobility distribution allows for more accurate predictions of disease transmission and enhances the precision of epidemiological models to reflect real-world scenarios. However, it is often impossible to measure the mobility of a whole population directly in real-world situations. In practice, the available data during epidemics is the number of infected individuals, which corresponds to the infected population ratio, defined as
@@
in the mobility-based SIRS model (4). In this section, we propose an inverse problem to infer the mobility distribution $f(m)\in\mathcal{L}^{1}(0,1)$ from the time-dependent infected ratio $Q_{0}^{I}(t)$. We prove that $f(m)$ is identifiable from $Q_{0}^{I}(t)$ (i.e., the inverse problem has a unique solution) and provide a machine-learning-based approach to infer $f(m)$ from real data.
3.1 Identifiability of the mobility distribution
In this section, we show that the observation $Q_{0}^{I}(t)$ contains sufficient information to infer the mobility distribution $f(m)$. In other words, this inference problem has a unique solution.